# BioMoDiffuse: Physics-Guided Biomechanical Diffusion for Controllable and Authentic Human Motion Synthesis

### v1

@@ Abstract
Human motion generation holds significant promise in fields such as animation, film production, and robotics. However, existing methods often fail to produce physically plausible movements that adhere to biomechanical principles. While recent autoregressive and diffusion models have improved visual quality, they frequently overlook essential biodynamic features, such as muscle activation patterns and joint coordination, leading to motions that either violate physical laws or lack controllability. This paper introduces BioMoDiffuse, a novel biomechanics-aware diffusion framework that addresses these limitations. It features three key innovations: (1) A lightweight biodynamic network that integrates muscle electromyography (EMG) signals and kinematic features with acceleration constraints, (2) A physics-guided diffusion process that incorporates real-time biomechanical verification via modified Euler-Lagrange equations, and (3) A decoupled control mechanism that allows independent regulation of motion speed and semantic context. We also propose a set of comprehensive evaluation protocols that combines traditional metrics (FID, R-precision, etc.) with new biomechanical criteria (smoothness, foot sliding, floating, etc.). Our approach bridges the gap between data-driven motion synthesis and biomechanical authenticity, establishing new benchmarks for physically accurate motion generation.
Figure 1: Our model, BioMoDiffuse, generates physically realistic motions based on textual input. By embedding Euler-Lagrange dynamics into the generation process, the model produces physical-optimized trajectories that inherently satisfy physical laws, achieving superior motion authenticity compared to current SOTA approaches. Specifically, BiMoDiffuse mitigates common kinematic artifacts including foot sliding, body interpenetration, and motion discontinuities while ensuring temporal smoothness.
1 Introduction
Human motion generation has wide-ranging applications in animation, film, and robotics. In recent years, there has been growing focus on generating human motion conditioned on language, sound, and motion trajectories [68, 61, 12, 49]. Among these, language-guided motion generation stands out due to its high practical value, complexity, and the extensive foundational research supporting the field. Recent advancements [15, 48, 66, 17, 22, 24, 50, 88] in autoregressive and diffusion models have significantly improved the synthesis of diverse motion sequences while maintaining semantic consistency.
@@ 1 Introduction
While existing methods have made strides in enhancing the visual plausibility of motion sequences, studies in biomechanics and animation highlight significant shortcomings[54, 10] of them in capturing key aspects of human motion. The absence of explicit modeling for biomechanical and dynamic features prevents current models from learning essential information, such as acceleration patterns and muscle activation levels. Consequently, generated motions are often rigid and physically unrealistic, deviating significantly from natural human movement in rhythm and joint coordination. Additionally, current evaluation metrics fail to quantitatively assess physical plausibility and motion velocity, while existing methods struggle with precise control over movement amplitude and speed. These limitations considerably diminish the practical value of motion generation in professional applications that require biomechanical accuracy.
@@
To address these challenges, we propose BioMoDiffuse, a biomechanical-aware motion generation framework that seamlessly integrates musculoskeletal dynamics with diffusion processes. Our key innovation lies in embedding explicit physical supervision into the diffusion pipeline, leveraging a physics-aware network based on modified Euler-Lagrange equations and physics guidance to enforce biomechanical plausibility. Specifically, BioMoDiffuse introduces real-time biomechanical verification during denoising and a decoupled motion control system for independent regulation of velocity and semantics. Additionally, since current evaluation protocols neglect biomechanical-related aspects, we establish a set of comprehensive evaluation metrics emphasizing physical realism, combining biomechanical validity metrics (e.g., motion smoothness, foot sliding rate) with traditional generation metrics (FID, Diversity). Experiments on HumanML3D [14] and KIT-ML [51] demonstrate significant improvements over existing methods.
The main contributions of this paper are as follows:
  • A biomechanics-guided diffusion framework that integrates EMG signals and kinematic features with acceleration-based physical supervision, ensuring compliance with rigid-body dynamics .
  • A physics-aware network architecture based on Euler-Lagrange principles, incorporating a novel motion control system for enhancing controllability.
  • A comprehensive evaluation framework that emphasizes physical realism, unifying biomechanical metrics and traditional metrics.
2 Related Work
Physical plausible motion processing. In motion generation, reinforcement learning (RL) has been utilized to fine-tune models for physical plausibility [72, 89, 67]. For example, ReinDiffuse [18] and PhysDiff [77] extend MDM [66] to improve motion quality. However, these RL-based approaches predominantly focus on refining foot-ground contact using opaque, black-box reward functions, often overlooking broader rigid-body dynamics and biomechanical principles. The lack of transparency in their loss design further constrains interpretability and limits precise control over kinematic behaviors. Alternatively, some methods incorporate auxiliary constraints during training [75, 29, 53, 34]; for instance, OmniControl [75] leverages spatial signals to guide generation. Nonetheless, these approaches tend to prioritize controllability at the expense of enforcing comprehensive physical laws, disproportionately emphasizing foot-centric adjustments rather than ensuring holistic motion plausibility.
Physical plausibility is also addressed in other motion-related tasks that demand biomechanically authentic synthesized motion sequences, such as human pose estimation from videos [32, 37, 33, 63, 60]. Recent advances try to incorporate physics-based modeling [78, 58, 87, 86, 59, 52, 35, 74] to enhance the quality of the estimated motion sequences. However, these methods still face challenges of suboptimal precision and excessive jitter. Although techniques like Gaussian filtering can suppress high-frequency noise, they often compromise precision. Consequently, several studies [76, 7, 25, 79, 69, 30, 65] have explored pose refinement techniques to reduce temporal instability. For instance, SmoothNet [79] employs a parametric approach to optimize the temporal dimension, substantially enhancing motion quality.
@@ 2 Related Work
Diffusion models. Diffusion models have demonstrated remarkable success in image generation, with various methods leveraging language guidance to produce high-quality images [55, 57, 81]. Among these approaches, DDPM [20] has emerged as the most widely adopted sampler, achieving impressive results. Building on DDPM, DDIM [62] reduces the number of inference steps while enabling deterministic sampling by adjusting the variance. The introduction of Classifier Guidance (CG) [9] empowered diffusion models to handle conditional generation tasks effectively, and CFG [19] further streamlined this process by replacing CG, enhancing the efficiency of guidance integration during training. In this paper, we observe that incorporating additional domain knowledge parallels the mechanisms of CG and CFG conceptually, leading us to re-derive the diffusion process accordingly.
@@
In recent years, many studies have tackled the Diffusion Inversion problem, which provides a method for mapping real images to latent space, with broad applications in image editing [46, 71, 36, 26, 85, 8]. However, Diffusion Inversion methods have rarely been applied to tasks outside of image generation. We observe that both motion refinement and image editing aim to preserve the macroscopic structure while modifying finer details. Consequently, we are the first to apply DDIM Inversion to the task of motion refinement.
@@
Text-driven human motion generation. Early approaches to text-driven human motion generation conceptualized motion as a form of language, adopting autoregressive methods commonly used in language models [2, 1, 48, 3]. Recognizing the potential of large language models, subsequent studies improved motion embeddings to enable the joint learning of motion and text, yielding promising results [15, 16, 80, 11, 73, 38, 42].
@@
MDM [66] pioneered the application of diffusion models to human motion generation, while MLD [5], inspired by VAE [31], introduced a diffusion process within the latent space. MotionDiffuse [82] established a larger model framework, using cross-attention to link motion and text. Building on MotionDiffuse, ReMoDiffuse [83] incorporated retrieval-based motion databases and supplementary knowledge, resulting in significant performance improvements. MoFusion [43] and StableMoFusion [22], drawing inspiration from Stable Diffusion [55], utilized the UNet [56] architecture without relying on external knowledge, achieving state-of-the-art results. Notably, experiments with StableMoFusion revealed that DiT [47] and RetNet [64] underperformed relative to UNet, validating our choice of UNet as the backbone for the diffusion framework.
3 Method
Current approaches fail to capture the essential characteristics of human motion due to the absence of biomechanical constraints and rigid-body dynamics modeling, limiting their ability to learn authentic movement patterns. To address this gap, we present a novel motion generation framework that explicitly incorporates biomechanical principles and rigid-body kinematics, which is capable of synthesizing authentic physically plausible human motions. Our methodological contributions are as follows: Sec. 3.1 establishes the theoretical foundation by formalizing the biomechanical constraints; Sec. 3.2 introduces a physics-aware network architecture derived from Euler-Lagrange equations, ensuring dynamic compliance; Sec. 3.3 presents our biomechanics-guided diffusion based on joint probability distribution analysis. The complete motion generation framework is detailed in Sec. 3.4. Furthermore, Sec. 3.5 demonstrates that our proposed method, equipped with diffusion inversion, can be effectively utilized for physics-aware motion refinement.
Figure 2: Overview of BioMoDiffuse. At each diffusion timestep $t$, our framework processes the motion sequence $\Theta^{t}$ alongside text conditions. The text is encoded using a frozen CLIP encoder and concatenated with $\Theta^{t}$ as input to the UNet backbone. The light brown panel (left) illustrates our physics-aware network, which consists of: (1) An Accel Net that implements augmented Euler-Lagrange equations to predict biomechanically valid accelerations, and (2) An Autoregressive Block that models temporal dependencies. The training objectives combine: $\mathcal{L}_{\text{phys}}$ (from Accel Net), $\mathcal{L}_{\text{diffusion}}$ (reconstruction), and $\mathcal{L}_{\text{accel}}$ (acceleration supervision against ground truth). Structural details of the orange (autoregressive block) and light blue (Accel Net) components are provided in the right panels.
3.1 Preliminaries
@@
Text-guided human motion generation requires an input text to condition the generation of a motion sequence $\Theta\in\mathbb{R}^{T\times D}$, where $T$ denotes the sequence length and $D$ represents the dimension of the joint representation. This joint representation can be converted into a coordinate representation $\Phi\in\mathbb{R}^{T\times 3d}$, where $d$ is the number of joints. The Euler-Lagrange equation is given by:
$F=M\cdot\ddot{\Phi_{t}}+C,$ (1)
@@
where $R$ is the muscle rotation matrix, $A(u)$ is the muscle activation force, and $u$ is the muscle activation coefficient. Using this decomposition, the Euler-Lagrange equation for human motion can be rewritten as Eq. 3. Please refer to Appendix D for details.
$\ddot{\Phi_{t}}=M^{-1}(E+C+R\cdot A(u)).$ (3)
3.2 Physics-Aware Network
The Euler-Lagrange equation-based network provides a stronger prior than conventional neural networks, as it inherently captures acceleration dynamics while reducing problem complexity. A crucial design here is to directly embed the constraints within the network architecture itself. In a physical sense, the Euler-Lagrange equation serves as a rigorous constraint for human motion: by framing motion generation as $Motion=Gen(text)$, this equation ensures the learned function aligns closely with real-world biomechanics; From a deep learning perspective, Euler-Lagrange equation networks offer a more robust alternative to traditional non-parametric constraints for handling diverse human skeleton data in supervised settings. Additionally, the inherent properties of Euler-Lagrange equations in our problem context guarantee convergence to a stable and unique solution.
@@ 3.2 Physics-Aware Network
Following the MIA dataset representation, a learnable mapping exists between human poses and the activation coefficients of primary muscles $u\in\mathbb{R}^{8}$, defined as $u=u(\Phi_{:t})$. The muscle rotation matrix $R$ depends solely on pose rotation, while the inertia matrix $M$—a symmetric matrix—is intrinsic to the body and determined by position. The Coriolis force and gravity are formulated as:
$C=\dot{M}(\Phi_{t})\dot{\Phi_{t}}-\frac{1}{2}\dot{\Phi_{t}}^{T}\frac{\partial M}{\partial\Phi_{t}}\dot{\Phi_{t}}+G(\Phi_{t}),$ (4)
@@
which are also pose-dependent. External contact forces $F$ (e.g., elastic/resistive forces) depend on both position and velocity. Thus, Eq. 3 is reformulated as:
$\ddot{\Phi}_{t}=M^{-1}(\Phi_{:t})\left(F(\Phi_{:t})+C(\Phi_{:t})+R(\Phi_{:t})\odot A(u)\right),$ (5)
where $\odot$ denotes element-wise multiplication. We implement three neural networks to approximate $M$, $F$, and $C$, with $M$ constrained as symmetric. The term $R\odot A(u)$ is modeled via cross-attention, enabling lightweight networks to predict pose acceleration. Finally, leveraging autoregressive processing, the predicted $\ddot{\Phi}_{t}$ and speed control signal are integrated as inputs to derive the optimized $\Theta_{t}$, ensuring temporal coherence and biomechanical fidelity.
3.3 Physics Guidance
We introduce additional guidance into the diffusion process to ensure that acceleration remains physically meaningful and achieve more stable training and faster convergence. Our network explicitly treats the acceleration $\ddot{\Phi_{t}}$ as an intermediate variable. We aim for the acceleration $\ddot{\Phi_{t}}$ in our physics-based network to converge to the real result as quickly as possible. To achieve this, we impose a supervised constraint on the acceleration during the diffusion process. Specifically, at each reverse diffusion step, we consider extending the distribution at step $t$, denoted by $\Theta^{t}$, into a joint distribution $S^{t}=\{\Theta^{t},\ddot{\Phi}^{t}\}$. According to DDPM theory, when performing diffusion on $S^{t}$, the model’s loss function can be written as:
$\min E_{Q(S^{0}:S^{T})}\left[-\log P_{\theta}(S^{0}|S^{1})\right.$ (6)
We argue that, during the denoising process, the quantities $\Theta^{t-1},\ddot{\Phi}^{t-1}$ at step $t-1$ are unknown and do not satisfy the interpolation relationship. However, the known quantities $\Theta^{t},\ddot{\Phi}^{t}$ at step $t$ do satisfy this relationship. This leads to the following:
$\log P_{\theta}(S^{t-1}|S^{t})=\log P_{\theta}(\Theta^{t-1},\ddot{\Phi}^{t-1}|\Theta^{t},\ddot{\Phi}^{t})$ (7)
$=\log P_{\theta}(\Theta^{t-1},\ddot{\Phi}^{t-1}|\Theta^{t})$
$=\log P_{\theta}(\Theta^{t-1}|\ddot{\Phi}^{t-1},\Theta^{t})+\log P_{\theta}(\ddot{\Phi}^{t-1}|\Theta^{t}).$
In the expression in Eq. 7, $\log P_{\theta}(\Theta^{t-1}|\ddot{\Phi}^{t-1},\Theta^{t})$ resembles the condition $c$ in CG and CFG, while $\log P_{\theta}(\ddot{\Phi}^{t-1}|\Theta^{t})$ can be obtained using the physical network we propose. Thus, during training, we omit the physical network with probability $P_{c}$, and use it with probability $1-P_{c}$. We set $\ddot{\Phi}^{t-1}=\text{PhysNet}(\Theta^{t})$, and the predicted acceleration is compared to the ground truth acceleration, introducing a physics loss $L_{\text{phys}}=\text{MSE}(\text{PhysNet}(\Theta^{t}),\ddot{\Phi_{\text{gt}}}^{t})$. To further enhance acceleration accuracy, we design the following additional loss term for supervision:
$L_{\text{accel}}=\text{MSE}(\ddot{\Phi}^{t},\ddot{\Phi_{\text{gt}}}^{t}).$ (8)
Thus, the total loss function during training can be written as:
$L=L_{\text{diffusion}}+L_{\text{accel}}+L_{\text{phys}}.$ (9)
3.4 Model Architecture
@@
Table 1: Quantitative evaluation on the HumanML3D and KIT-ML test set. $\pm$±indicates a 95% confidence interval. Bold face indicates the best result, while underscore refers to the second best. We reimplemented StableMoFusion based on the official code.
@@
Our biomechanics-aware diffusion framework is built upon three core components, as shown in Fig. 2. The architecture seamlessly integrates biomechanical priors with diffusion dynamics through:
Biomechanical Encoding. Using the MIA dataset [6], we pre-train a neuromuscular encoder that establishes a dynamical mapping $\mathcal{F}:\Theta\rightarrow u$ between joint angles $\Theta$ and muscle activation coefficients $u$. This encoder remains differentiable throughout end-to-end training.
Physics-Conditioned Diffusion. The UNet-based diffusion backbone processes motion sequences $\mathbf{x}_{t}$ at timestep $t$ through:
$z_{\theta}=\text{BioMoDiffuse}\big{(}\text{CLIP}(\mathbf{c}),\mathbf{x}_{t},\mathcal{F}(\Theta)\big{)},$ (10)
where text conditioning $\mathbf{c}$ is incorporated via cross-attention layers. We adopt a two-phase training process: initial diffusion followed by physics-constrained fine-tuning (which introduces the physics loss $\mathcal{L}_{\text{phys}}$).
Controllable Generation. Our decoupled control mechanism enables independent manipulation of motion dynamics and semantics through: (1) classifier-free guidance and physics guidance weights, and (2) the injection of biomechanical constraints via the pre-trained encoder.
3.5 Pose Refinement with Diffusion Inversion

### v2

# Biomechanics-Guided Residual Approach to Generalizable Human Motion Generation and Estimation

@@ Abstract
Human pose, action, and motion generation are critical for applications in digital humans, character animation, and humanoid robotics. However, many existing methods struggle to produce physically plausible movements that are consistent with biomechanical principles. Although recent autoregressive and diffusion models deliver impressive visual quality, they often neglect key biodynamic features and fail to ensure physically realistic motions. Reinforcement Learning (RL) approaches can address these shortcomings but are highly dependent on simulation environments, limiting their generalizability. To overcome these challenges, we propose BioVAE, a biomechanics-aware framework with three core innovations: (1) integration of muscle electromyography (EMG) signals and kinematic features with acceleration constraints to enable physically plausible motion without simulations; (2) seamless coupling with diffusion models for stable end-to-end training; and (3) biomechanical priors that promote strong generalization across diverse motion generation and estimation tasks. Extensive experiments demonstrate that BioVAE achieves state-of-the-art performance on multiple benchmarks, bridging the gap between data-driven motion synthesis and biomechanical authenticity while setting new standards for physically accurate motion generation and pose estimation.
Index Terms:
Human Motion Generation, Pose Estimation, Diffusion Models.
(a) Results across multiple tasks
(b) Quantitative comparison
Figure 1: (a) Demonstrates our method’s results across multiple tasks, including motion generation, motion refinement, pose estimation, and motion editing. (b) Shows a quantitative comparison of different metrics against current state-of-the-art methods. Our proposed BioVAE framework demonstrates strong generalization across these tasks. Compared to state-of-the-art methods, our approach generates motions with significantly reduced foot floating, penetration, and clipping artifacts, as well as improved temporal smoothness. Notably, BioVAE achieves a clear improvement in physical plausibility metrics, setting a new benchmark for biomechanically realistic motion synthesis.
footnotetext: Zixi Kang, Xinghan Wang and Yadong Mu are with the Wangxuan Institute of Computer Technology, Peking University, Beijing 100871, China (e-mail: forever.kzx0713@stu.pku.edu.cn, xinghan_wang@pku.edu.cn, myd@pku.edu.cn). Yadong Mu is the corresponding author.
I Introduction
Human pose, action, and motion generation have wide-ranging applications in motion capture, animation, and robotics. In recent years, there has been a growing focus on generating human motion conditioned on image, language, sound, and motion trajectories [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
@@ I Introduction
However, studies in biomechanics and animation[13, 14] have shown that current methods fail to capture the distinct characteristics of human motion, resulting in generated sequences that often deviate significantly from natural movement, with unnatural rhythms and unrealistic joint configurations. Moreover, the lack of explicit modeling of biodynamic features prevents these models from capturing critical information such as acceleration patterns and muscle activation levels, often resulting in motions that appear stiff or violate physical laws. Furthermore, existing evaluation metrics also overlook the quantitative assessment of physical plausibility and motion velocity, while current approaches struggle to control movement amplitude and speed precisely. These limitations severely undermine the practical value of motion generation in professional applications that demand high biomechanical precision.
@@
To address these challenges, we propose BioVAE, a biomechanics-aware framework that seamlessly integrates musculoskeletal dynamics into diffusion processes. Our key innovation lies in embedding Euler-Lagrange equations and EMG signals into the VAE architecture, while connecting the VAE residuals to the diffusion process to enable end-to-end, generalizable motion generation. Furthermore, since current evaluation protocols largely overlook biomechanical aspects, we establish a comprehensive set of metrics which emphasizes physical realism, combining biomechanical validity measures (e.g., motion smoothness, foot sliding rate) with traditional generation metrics such as FID and diversity. Extensive experiments on HumanML3D[15], KIT-ML[16], 3DPW[17], and RICH[18] demonstrate that BioVAE achieves significant improvements over existing methods.
The main contributions of this paper are as follows.
  • We propose BioVAE, a residual approach based on Euler-Lagrange equations that can be jointly trained with diffusion models in a self-supervised setting.
  • We introduce EMG signals and acceleration-based physical supervision, leveraging additional biomechanical priors to enhance the model’s capability.
  • We design extensive experiments across different tasks to evaluate both reconstruction quality and physical plausibility, demonstrating the strong generalizability of our framework.
II Related Work
II-A Physically plausible motion processing.
In motion generation, Reinforcement Learning (RL) has been utilized to fine-tune models for physical plausibility [19, 20, 21]. For example, ReinDiffuse [22] and PhysDiff [23] extend MDM [24] to improve motion quality. However, these RL-based methods suffer from two major limitations: (1) they lack the ability of generalization and require task-specific, dataset-specific, and model-specific reward designs; and (2) they depend heavily on simulation environments, making the training process a black box and preventing end-to-end optimization.
Alternatively, some methods incorporate auxiliary constraints during training [25, 26, 27]; for instance, OmniControl [25] leverages spatial signals to guide generation. However, since these approaches mainly focus on enabling multimodal and editable motion generation instead of improving motion quality, they remain prone to challenges like limited precision and excessive jitter. Although techniques like Gaussian filtering can suppress high-frequency noise, they often degrade motion precision.
Consequently, several studies [28, 29, 30, 31, 32, 33, 34, 35] have explored pose refinement techniques to reduce temporal instability. For example, SmoothNet [31] introduces a parametric approach to optimize temporal coherence and substantially improve motion quality. Nevertheless, these methods do not address generalization and cannot be seamlessly integrated with diffusion models or large language models (LLMs), failing to leverage the powerful generative capabilities of such backbones and exhibiting inherent limitations.
II-B Diffusion models.
@@ II Related Work
Diffusion models have achieved remarkable success in image generation, with numerous approaches leveraging language guidance to produce high-quality results [36, 37, 38]. Among them, Denoising Diffusion Probablistic Models (DDPM) [39] has become the most widely adopted sampling method, delivering impressive results. Based on DDPM, DDIM [40] reduces inference steps and allows deterministic sampling by adjusting the variance schedule. The introduction of Classifier Guidance (CG)[41] extended diffusion models to conditional generation tasks, while Classifier-Free Guidance (CFG)[42] further streamlined this process by replacing CG and improving the efficiency of the guidance during training. Inspired by these advancements, we observe that incorporating domain-specific knowledge is conceptually aligned with the mechanisms of CG and CFG, thereby motivating a re-derivation of the diffusion process.
@@
In parallel, recent studies have addressed the diffusion inversion problem, which map real images to latent space and supports various image editing applications [43, 44, 45, 46, 47, 48]. However, diffusion inversion techniques have rarely been explored beyond image generation. This work is the first to apply DDIM inversion to motion generation tasks.
II-C Text-driven human motion generation.
@@
Early approaches to text-driven human motion generation conceptualized motion as a form of language, adopting autoregressive methods commonly used in language models [49, 50, 51, 52]. Recognizing the potential of large language models, subsequent studies improved motion embeddings to enable joint learning of motion and text, yielding promising results [53, 54, 55, 56, 57, 58].
@@
MDM [24] pioneered the application of diffusion models to human motion generation, while MLD [59], inspired by VAE [60], introduced a diffusion process within the latent space. MotionDiffuse [61] established a larger model framework, using cross-attention to link motion and text. Building on MotionDiffuse, ReMoDiffuse [62] incorporated retrieval-based motion databases and supplementary knowledge, resulting in significant performance improvements. MoFusion [63] and StableMoFusion [64], drawing inspiration from Stable Diffusion [36], utilized the UNet [65] architecture without relying on external knowledge, achieving state-of-the-art results. To enable our BioVAE to deeply integrate with the diffusion process, we formulate BioVAE as a self-supervised residual connection, re-derive the probability formulation of the diffusion process, and redesign the loss function to achieve stable end-to-end training of the two components.
III Method
Figure 2: Overview of BioVAE. We employ the MIA dataset [66] and adopt a training strategy that learns the mapping from motion sequences to EMG signals within the motion encoder. In the latent space, a neural network predicts four intermediate variables: $M^{-1}$, $E$, $C$, and $A$, which are then combined through the Euler-Lagrange equations to reconstruct the joint local rotational accelerations. The motion decoder iteratively generates residuals from the acceleration information to guide the UNet’s generation process. BioVAE can be applied to a variety of tasks and modalities, such as text-to-motion, pose estimation, motion editing, and motion refinement, demonstrating strong generalization across tasks while significantly enhancing the physical plausibility of the generated motions.
Existing methods fail to capture the unique characteristics of human motion due to the absence of biomechanical constraints and the lack of rigid body modeling. In this section, we present our BioVAE framework, which integrates biomechanical principles and rigid body dynamics into the network architecture, providing a diffusion-based model that achieves state-of-the-art performance in multiple tasks. We begin in Section III-A by introducing the preliminaries. Section III-B details the architecture of the BioVAE network. In Section III-C, we derive the joint distribution formulation that allows end-to-end training of BioVAE with diffusion through residual connections. Finally, Section III-D describes our diffusion inversion strategy, showing how BioVAE supports motion refinement, motion editing, and pose estimation.
III-A Preliminaries
@@
Human motion generation produces a motion sequence $\Theta\in\mathbb{R}^{T\times D}$, where $T$ denotes the length of the sequence and $D$ is the dimension of the joint representation. This motion sequence can be transformed into a joint coordinate representation $X\in\mathbb{R}^{T\times 3d}$, where $d$ represents the number of joint nodes. By modeling the human body as an articulated rigid-body system, the Euler-Lagrange equation can be formulated as:
$F=M\cdot\ddot{X_{t}}+C,$ (1)
@@
where $R$ is the muscle rotation matrix, $A(u)$ is the muscle activation force, and $u$ is the muscle activation coefficient. Using this decomposition, the Euler-Lagrange equation for human motion can be rewritten as Equation 3.
$\ddot{X_{t}}=M^{-1}(E+C+R\cdot A(u)).$ (3)
III-B BioVAE Architecture
Compared to conventional neural networks, a network based on the Euler-Lagrange equation provides a stronger prior, as it can inherently capture acceleration dynamics while leveraging constraints to reduce the complexity of the problem. In this section, we introduce how to design a VAE architecture guided by the Euler-Lagrange equation.
@@ III-B BioVAE Architecture
Following the representation in the MIA dataset, there exists a learnable mapping between human poses and the activation coefficients of the primary muscles $u\in\mathbb{R}^{8}$, defined as $u=u(X_{:t})$. The muscle rotation matrix $R$ depends solely on the rotational component of the pose, while the inertia matrix $M$ is an intrinsic property of the body determined only by position. The Coriolis force and gravity can be expressed as follows.
$C=m\omega^{2}r+2m\omega\times\dot{X}_{t}-\nabla U_{gravity}(X_{t}),$ (4)
@@
which are likewise pose-dependent. The external contact force $F$, including the elastic and resistive forces, depends on both position and velocity. Using this formulation, Equation (3) can be rewritten as a function of $X_{:t}$ in Equation (5):
$\ddot{X}_{t}=M^{-1}(X_{:t})\left(F(X_{:t})+C(X_{:t})+R(X_{:t})\cdot A(u)\right).$ (5)
We implement three neural networks to approximate $M^{-1}$, $F$, and $C$, where $M^{-1}$ is restricted to be symmetric. The term $R\cdot A(u)$ is modeled using cross-attention, allowing four lightweight networks to predict the acceleration of the pose effectively. Once acceleration $\ddot{X}_{t}$ is predicted, we employ an autoregressive decoder to reconstruct the residual between predicted $X_{t}$ and the ground truth. This enables optimization at each training time step while ensuring temporal consistency and biomechanical fidelity.
@@
TABLE I: Quantitative evaluation on the HumanML3D and KIT-ML test set. The confidence interval is indicated by±$\pm$. Bold face indicates the best result, while underscore refers to the second best. We reimplemented StableMoFusion based on the official code.
@@
III-C End-to-End Residual Training
In previous approaches, optimization methods such as reinforcement learning (RL) or refinement [67, 31] are often applied as post-processing on top of the backbone model. In contrast, we integrate BioVAE with diffusion to enable end-to-end training. Furthermore, we apply optimization at every diffusion training timestep, allowing our model to perform iterative refinement during inference, which substantially improves the quality of generation.
Traditional RL and simulation-based methods [21, 20, 23] typically require a complete human motion sequence and thus cannot be integrated into the noisy training process of diffusion models. To address this, our network explicitly treats acceleration $\ddot{X_{t}}$ as an intermediate variable. We introduce supervised constraints on the acceleration during the diffusion process to ensure that the predicted accelerations are physically meaningful, resulting in a more stable training process and faster convergence.
Specifically, at each reverse diffusion step, we consider extending the distribution at step $t$, denoted as $\Theta^{t}$, to a joint distribution $S^{t}=\{\Theta^{t},\ddot{X}^{t}\}$. According to DDPM theory, when performing diffusion on $S^{t}$, the model’s loss function can be written as:
$\min E_{Q(S^{0}:S^{T})}\left[-\log P_{\theta}(S^{0}|S^{1})\right.$ (6)
We argue that, during the denoising process, the quantities $\Theta^{t-1},\ddot{X}^{t-1}$ at step $t-1$ are unknown and do not satisfy the interpolation relationship. However, the known quantities $\Theta^{t},\ddot{X}^{t}$ at step $t$ do satisfy this relationship. This leads to the following:
$\log P_{\theta}(S^{t-1}|S^{t})=\log P_{\theta}(\Theta^{t-1},\ddot{X}^{t-1}|\Theta^{t},\ddot{X}^{t})$ (7)
$=\log P_{\theta}(\Theta^{t-1},\ddot{X}^{t-1}|\Theta^{t})$
$=\log P_{\theta}(\Theta^{t-1}|\ddot{X}^{t-1},\Theta^{t})+\log P_{\theta}(\ddot{X}^{t-1}|\Theta^{t}).$
In the expression in Equation 7, $\log P_{\theta}(\Theta^{t-1}|\ddot{X}^{t-1},\Theta^{t})$ resembles the condition $c$ in CG and CFG, while $\log P_{\theta}(\ddot{X}^{t-1}|\Theta^{t})$ can be obtained using the VAE architecture we propose. Thus, during training, we set $\ddot{X}^{t-1}=\text{VAE}(\Theta^{t})$, and the predicted acceleration is compared to the ground truth acceleration, introducing a physics loss $L_{\text{phys}}=\text{MSE}(\text{VAE}(\Theta^{t}),\ddot{X_{\text{gt}}}^{t})$. To further enhance acceleration accuracy, we design the following additional loss term:
$L_{\text{accel}}=\text{MSE}(\ddot{X}^{t},\ddot{X_{\text{gt}}}^{t}).$ (8)
Thus, the total loss function during training is written as:
$L=L_{\text{diffusion}}+L_{\text{accel}}+L_{\text{phys}}.$ (9)
III-D Generation with Diffusion Inversion